OpenStudy (anonymous):

An investor invested a total of $1300 in two mutual funds. One fund earned a 6% profit while the other earned a 4% profit. If the investor's total profit was $56, how much was invested in each mutual fund? the amount invested that earned 6% is? the amount invested that earned 4% is?

OpenStudy (anonymous):

you have a system of eqns, trick is always setting the thing up x + y = 1300 0.06x + 0.04y = 56 now you can solve using whichever method you want, prob substitution of one into the other will allow you to solve it the simplest and quickest...

OpenStudy (anonymous):

this is as far as i gotten the rest of the formula is confusing me

OpenStudy (anonymous):

Interest on both investments 6.00% x + 4.00% y= 56 Multiply by 100 6 x + 4 y= 5600.00

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

solve that for y then sub back into x + y =1300 y= 1400 - (3/2)x so x + (1400 - (3/2)x) = 1300 ---> x=200 therefore y =?

OpenStudy (anonymous):

hope that helps! :)

OpenStudy (anonymous):

you really don't need two variables. if the amount invested at 6% is \(x\) then the amount invested at 4% must be \(1300-x\) since the total was $1300

OpenStudy (anonymous):

you can go right to \[.06x+.04(1300-x)=56\] or better still \[6x+4(1300-x)=5600\]

OpenStudy (anonymous):

you good from there?